# Avoiding Premature Collapse: Adaptive Annealing for Entropy-Regularized Structural Inference

### v2

@@ Abstract
Differentiable matching layers, often implemented via entropy-regularized Optimal Transport, serve as a critical approximate inference mechanism in structural prediction. However, recovering discrete permutations via annealing $\epsilon\to 0$ is notoriously unstable. We identify a fundamental mechanism for this failure: Premature Mode Collapse. By analyzing the non-normal dynamics of the Sinkhorn fixed-point map, we reveal a theoretical thermodynamic speed limit. Under standard exponential cooling, the shift in the target posterior ($O(1)$) outpaces the contraction rate of the inference operator, which degrades as $O(1/\epsilon)$. This mismatch inevitably forces the inference trajectory into spurious local basins. To address this, we propose Efficient EPH-ASC, an adaptive scheduling algorithm that monitors the stability of the inference process. By enforcing a linear stability law, we decouple expensive spectral diagnostics from the training loop, reducing overhead from $O(N^{3})$ to amortized $O(1)$. Our implementation and interactive demo are available at https://github.com/xxx0438/torch-sinkhorn-asc and https://huggingface.co/spaces/leon0923/torch-sinkhorn-asc-demo. bounded away from zero in generic training dynamics unless the feature extractor converges unrealistically fast.
@@ 1 Introduction
Entropy-regularized Optimal Transport (OT) has become a standard surrogate for combinatorial inference in differentiable matching and sorting layers Peyré and Cuturi (2019). Practitioners often attempt to recover hard assignments (Maximum A Posteriori estimates) by annealing the regularization parameter $\epsilon\to 0$. Empirically, however, this cooling process is fragile: annealing frequently triggers premature sparsification, where the transport plan collapses onto spurious permutation vertices before the true structural signal has matured.
Existing analyses mostly treat this as a convergence nuisance or address it via static asymptotics. In contrast, we demonstrate that this failure is not merely an empirical misfortune, but a theoretical inevitability of standard exponential annealing. We reframe the failure as a sequential inference tracking problem: the algorithm must track a moving fixed point $P_{\epsilon_{t}}^{*}$. We prove that exponential schedules fundamentally violate the thermodynamic speed limit of the solver, guaranteeing that the instantaneous distributional shift will eventually exceed the solver’s restoring capacity.
@@ 2.2 Thermodynamic Sensitivity ($1/\varepsilon$)
To quantify the "velocity" of the distribution shift, we adopt a localized non-degeneracy assumption that fixes an active support set $S$ for small $\varepsilon$. Under this assumption, the sensitivity of the Sinkhorn map with respect to $\varepsilon$ scales like $1/\varepsilon$. This follows from implicit differentiation of the entropic optimality conditions Cominetti and Martín (1994); Weed (2018).

### v3

@@ Abstract
Differentiable matching layers and residual connection paradigms, often implemented via entropy-regularized Optimal Transport (OT) Cuturi (2013); Peyré and Cuturi (2019), serve as critical mechanisms in structural prediction and architectural scaling. However, recovering discrete permutations or maintaining identity mappings via annealing $\epsilon\rightarrow 0$ is notoriously unstable Peyré and Cuturi (2019). We identify a fundamental mechanism for this failure: Premature Mode Collapse. By analyzing the non-normal dynamics of the Sinkhorn fixed-point map Trefethen and Embree (2005), we reveal a theoretical thermodynamic speed limit: standard exponential cooling outpaces the contraction rate of the inference operator, which degrades as $O(1/\epsilon)$ Cominetti and San Martín (1994); Weed (2018). To address this, we propose Efficient Piecewise Hybrid Adaptive Stability Control (EPH-ASC), an adaptive scheduling algorithm that monitors the stability of the inference process. We demonstrate that EPH-ASC is essential for stabilizing Manifold-Constrained Hyper-Connections (mHC) DeepSeek-AI et al. (2025) during large-scale training on the FineWeb-Edu dataset, preventing late-stage gradient explosions by enforcing a linear stability law.
@@ 1 Introduction
Entropy-regularized Optimal Transport (OT) has become a standard surrogate for combinatorial inference Cuturi (2013); Altschuler et al. (2017) and modern macro-architecture design DeepSeek-AI et al. (2025); Zhu et al. (2024). Practitioners often attempt to recover hard assignments or specialized routing by annealing the regularization parameter $\epsilon\rightarrow 0$ Genevay et al. (2018); Peyré and Cuturi (2019). Recently, studies such as Manifold-Constrained Hyper-Connections (mHC) DeepSeek-AI et al. (2025) have utilized the Sinkhorn-Knopp algorithm Sinkhorn and Knopp (1967) to project residual mappings onto the Birkhoff polytope, thereby restoring the identity mapping property in multi-stream architectures DeepSeek-AI et al. (2025).
Empirically, however, this cooling process is fragile Peyré and Cuturi (2019); Blondel et al. (2018). As $\epsilon$ decreases, the sensitivity of the optimal plan to cost perturbations blows up as $O(1/\epsilon)$ Cominetti and San Martín (1994); Weed (2018), a phenomenon we define as the Thermodynamic Speed Limit. In the presence of high-variance gradients from real-world datasets like FineWeb-Edu, standard exponential annealing Chizat (2024) violates this limit, causing the composite mapping in HC architectures to diverge or collapse DeepSeek-AI et al. (2025). We propose EPH-ASC to reconcile this instability by monitoring the Primal Drift Weed (2018); Eisenberger et al. (2022) and triggering a "Thermodynamic Pause" when the distributional shift exceeds the solver’s restoring capacity.
@@ 2.2 Thermodynamic Sensitivity ($1/\varepsilon$)
To quantify the "velocity" of the distribution shift, we adopt a localized non-degeneracy assumption that fixes an active support set $S$ for small $\varepsilon$. Under this assumption, the sensitivity of the Sinkhorn map with respect to $\varepsilon$ scales like $1/\varepsilon$. This follows from implicit differentiation of the entropic optimality conditions Cominetti and San Martín (1994); Weed (2018).
@@ 5.2 Results: Entropy and Convergence
5.3 Thermodynamic Robustness in Large-Scale LM Training
To rigorously assess the robustness of EPH-ASC under real-world conditions characterized by high gradient variance, we scale the experiment to a language modeling task using the FineWeb-Edu dataset[cite: 536].
Setup. We employ a lightweight NanoGemma architecture equipped with Manifold-Constrained Hyper-Connections (mHC). Unlike controlled benchmarks, this experiment utilizes a GPT-2 tokenizer and natural language data, introducing heavy-tailed noise into the optimization landscape[cite: 538]. We compare a standard exponential schedule (Naive) against EPH-ASC over 1,000 steps[cite: 539].
Late-Stage Collapse. As shown in Fig. 5 (Left), the Naive schedule (Red) appears successful for 98% of the trajectory but suffers a catastrophic gradient explosion at Step 980[cite: 541, 542]. This confirms that Sinkhorn instability is a latent risk that manifests when $\epsilon$ breaches the operator’s condition number limit.
Adaptive Intervention. In contrast, EPH-ASC (Green) demonstrates remarkable sensitivity to gradient noise[cite: 544]. Despite the jagged loss landscape, the controller detects critical distributional drift at Step 640—long before visible loss degradation. By triggering "Thermodynamic Braking" (Orange markers), it locks temperature at $\epsilon\approx 0.04$, securing a 340-step safety margin[cite: 546, 547]. The entropy plot (Fig. 5, Right) further reveals that EPH-ASC maintains a stable low-entropy regime, preventing the numerical underflow observed in the baseline[cite: 548].
Figure 5: Verification on FineWeb-Edu. Left: Real-world training loss shows deceptive stability followed by sudden Naive collapse[cite: 554]. Center: EPH-ASC detects instability at Step 640, creating a massive safety margin via braking[cite: 555]. Right: Entropy preservation prevents numerical underflow in the feature maturity phase[cite: 556].